# CantonMT: Cantonese to English NMT Platform with Fine-Tuned Models using Synthetic Back-Translation Data

### v1

@@ 3.3 Adding more real data
In the extension of our work, we managed to fine-tune the baseline models using more real data from another source called Wenlin https://wenlin.com/ where we obtained another 14.5K parallel Cantonese English dictionary corpus. We are curious about the model performances using more real data in addition to the 38K training corpus from words.hk. We listed the comparison scores in the second cluster of Figure 1 where it shows that the newly fine-tuned NLLB-200 using 52.5K data (38+14.5K=52.5K) produced higher scores on all metrics in comparison to 38K trained model; mBART fine-tuned using 52.5K gets better scores on three metrics except for SacreBLEU; Opus-MT surprisingly did not get any increase across metrics. Nevertheless, these outcomes demonstrated the possibility of improving model performances with more available real data, at least for NLLB and mBART models. Furthermore, data quality matters: simply adding 14.5K real data NLLB fine-tuned produced higher scores (underlined scores) than the best synthetic system that used 38x2=76K data.
@@ 4.1 User Interface
To test out the User Interface and different models for translation, users can just choose different model types and source languages, which dynamically capture the available model from the server, and allow users to select different training methods for the model. One can then type the sentence in the input box and click the translate button for the translation output from the model. The application layout is quite modular in case different model types or languages are added to the system, which could potentially be used as a base framework for different translation systems and simply add more languages to the input and output if one wishes to expand the implementations. This web application has taken a template (Wrigley, 2023) for an AI Code translator and modified it to fit the need, which is developed in TypeScript with the Next.js framework. The reason for choosing this framework is that it provides a very modern and minimalistic approach to web development.
@@ 4.2 Server
A Diagram outlining the modules can be seen in Figure 5 to understand the general structure of the server. Users can easily run the server on their local machines and it is well-documented in the Readme file in the open-source platform. The server has two main functionalities, where the first one will output the list of model paths given the model type and source languages. The second one provides the translation, where one could provide the details of the model and also the sentence in the language specified, and the server would respond with the translated sentence using the model output.
During the implementation, due to a lack of memory space, the server crashed multiple times on our local machine. To account for the potential crash, a model manager was produced, which implements an LRU cache for the different model loaders, where the least recently used model will be deleted from memory if it exceeds the limit of the number of models. The server is built entirely on the Python Flask library. The reason for choosing this framework is that the models can be run on the Python Transformers library, which provides a seamless implementation without much additional effort.
@@ 5 Discussion and Conclusion
In this work, we investigated the back-translation methodology for bilingual synthetic data generation for the sake of data augmentation for NMT, on a new language translation direction Cantonese to English. We tested both smaller-sized OpusMT and extra-large LLMs NLLB and mBART both using available bilingual real data and the larger synthetic data. The experiments show that all the fine-tuned models outperformed the baseline deployment models with large margins. Furthermore, the synthetic model nllb-syn-1:1-mbart produced higher scores using model switch method compared to without model switch. Last but not the least, the best performing fine-tuned models have similar (or even higher) evaluation scores than the current translators from IT companies of Baidu and Microsoft-Bing, although the fine-tuned GPT4 has the highest ones.
In the concern of data privacy such as sensitive data (e.g. clinical domain Han et al. (2024)), CantonMT can be fully controlled by users without interference from any third parties. We open-source our platform so that researchers can continue to integrate new models into the toolkit to promote Cantonese-English MT. We also plan to carry out human evaluations on the outputs from different systems to get more insights into the system errors.
@@ Acknowledgements
We thank words.hk and wenlin.com for the data usage. LH and GN thank the support from Grant “Assembling the Data Jigsaw: Powering Robust Research on the Causes, Determinants and Outcomes of MSK Disease” (The project has been funded by the Nuffield Foundation, but the views expressed are those of the authors and not necessarily the Foundation. Visit www.nuffieldfoundation.org) and Grant EP/V047949/1 “Integrating hospital outpatient letters into the healthcare data space” (funder: UKRI/EPSRC).
@@ Appendix A Appendix
Example raw text extracted from LIHKG website can be seen in Figure 6 before cleaning. The model parameters from OpusMT, extra-large NLLB and mBART are shown in Figure 7, which shows that NLLB and mBART have doubled the number of Transformer layers and almost 10 times more parameters than OpusMT.

### v3

@@ 3.3 Adding More Real Data
In the extension of our work, we managed to fine-tune the baseline models using more real data from another source called Wenlin [^15] https://wenlin.com/ where we obtained another 14.5K parallel Cantonese-English dictionary. We are curious about the model performance using more real data in addition to the 38K training corpus from words.hk. We listed the comparison scores in the second cluster of Table 3 where it shows that the newly fine-tuned NLLB-200 using 52.5K data (38+14.5K=52.5K) produced higher scores on all metrics in comparison to 38K trained model; mBART fine-tuned using 52.5K obtains better scores on three metrics except for SacreBLEU; Opus-MT surprisingly did not get any increase across the metrics. Nevertheless, these outcomes demonstrated the possibility of improving model performance with more available real data, at least for the NLLB and mBART models. Moreover, data quality matters: simply adding 14.5K real data to fine-tune NLLB produced higher scores (underlined scores) than the best synthetic system that used 38x2=76K data. Subsequently, when we managed to get another 10K real data from Opus corpus, it shows continuous improvement by training using all three corpus we have, located in the Cluster 4 bottom of the table.
@@ 4.1 User Interface
To test the user interface and different models for translation, users can choose from different model types and source languages, which dynamically capture the available models in the server, and allow users to select different training methods for the model. One can then type the source sentence in the input box and click the “Translate” button to obtain the translation output from the model. The application layout is quite modular in case different model types or languages are added to the system, which could potentially be used as a base framework for different translation systems. It is possible to simply add more languages to the input and output if one wishes to expand the implementations. The look-and-feel of this web application is based on a template (Wrigley, 2023) for an AI Code translator, which was customised and developed in TypeScript with the Next.js framework. The reason for choosing this framework is that it provides a very modern and minimalistic approach.
@@ 4.2 Server
A diagram outlining the modules can be seen in Figure 3 to understand the general structure of the server. Users can easily run the server on their local machines by following the instructions provided in a README file. The server has two main functionalities, where the first one will output the list of model paths given the model type and source languages. The second one provides the translation, where one could provide the details of the model and also the sentence in the language specified, and the server would respond with the translated sentence using the model output.
During our implementation, due to memory constraints, the server crashed multiple times on our local machine. To mitigate the risk of server crashes, a model manager was produced, which implements a Least Recently Used (LRU) cache for the different model loaders, where the least recently used model will be deleted from memory if it exceeds the limit of the number of models. The server is built entirely based upon the Python Flask library. The reason for choosing this framework is that the models can be run on top of the Python Transformers library, which provides seamless implementation without much additional effort.
@@ 5 Discussion and Conclusion
In this work, we investigated the back-translation methodology for bilingual synthetic data generation for the sake of data augmentation for NMT, on a new language translation direction, Cantonese-to-English. We tested both smaller-sized OpusMT and extra-large LLMs NLLB and mBART both using available bilingual real data and larger synthetic data. Our experiments show that all the fine-tuned models outperformed the baseline deployment models with large margins. Furthermore, the synthetic model nllb-syn-1:1-mbart produced higher scores using the model switch method compared to those without the model switch. Lastly, the best performing fine-tuned models have similar (or even higher) evaluation scores than the current commercially available translators of Baidu and Microsoft-Bing.
In terms of concerns of data privacy such as handling of sensitive data (e.g., in clinical applications related to health analytics of patient data Han et al. (2024)), CantonMT can be fully controlled by users without interference from any third parties. We open-source our platform so that researchers can continue to integrate new models into the toolkit to promote Cantonese-English MT. We also plan to carry out human evaluations on the outputs from different systems to get more insights into the system errors.
Limitations
The synthetic data generated in this work is based on the fine-tuned model using 38K words.hk bilingual dictionary corpus, the first corpus we managed to find. This restricted the synthetic data quality. In the following-up work, we plan to use the further fine-tuned model on all three corpora, words.hk, wenlin, and opus-10K, to generate better back-translated synthetic data. We expect this will improve the synthetic data fine-tuned models.
The whole procedure of how difficult it was to collect real Cantonese-English bilingual data shows that Cantonese-English MT is still at its beginning stage with many obstacles and challenges to public research.
@@ Acknowledgements
We thank words.hk and wenlin.com for the data. LH and GN thank the support from the following grants: (1) “Assembling the Data Jigsaw: Powering Robust Research on the Causes, Determinants and Outcomes of MSK Disease”. This project has been funded by the Nuffield Foundation (www.nuffieldfoundation.org), but the views expressed are those of the authors and not necessarily of the Foundation; and (2) the UKRI/EPSRC Grant EP/V047949/1 “Integrating hospital outpatient letters into the healthcare data space”.
@@ Appendix A Appendix
Example raw text extracted from LIHKG website can be seen in Figure 5 before cleaning. CantonMT user-friendly interface is shown in Figure 4 (Frontend: TypeScript with Next.js. Backend: Python - Flask).
The model parameters from OpusMT, extra-large NLLB and mBART are shown in Table 2, which shows that NLLB and mBART have doubled the number of transformer layers and have almost 10 times more parameters than OpusMT.
Opus NLLB mBart
Layers 12 24 24
Hidden Unit 512 1024 1024
Model Parameters 77.9M 615M 610.9M
Language Pair No Yes No
Release Year 2020 2022 2020
Table 2: Parameters from deployed models. Language pair: if the model contains Cantonese-English as a language pair
Explanation of Abbriviations used in the scoring table:
  • “nllb-forward-bl”: NLLB fine-tuned model in the forward translation direction (Cantonese-English) using the real 38K bilingual corpus
  • “nllb-forward-syn-h:h”: NLLB fine-tuned model using forward-translation generated synthetic data to substitute half of the 38K real data, i.e. 19K real and 19K synthetic
  • “nllb-forward-syn-1:1”: NLLB fine-tuned model using forward-translation generated synthetic data with the ratio 1:1, i.e. 38K real and 38K synthetic
  • “nllb-forward-syn-1:1-10E”: the same with above corpus setting but running 10 epochs, default is 3 epochs only
  • “nllb-forward-syn-1:1-mbart”: NLLB model fine-tuning using forward-translation generated synthetic data by another model mBART, 38K real and 38K synthetic
Model Name SacreBLEU BERTscore COMET
mBART-back-bl 20.3841 0.7944 0.8095
mBART-back-syn-1:1-NLLB+ 20.1923 0.7921 0.8068
nllb-back-bl 18.4713 0.7877 0.7927
nllb-back-syn-1:1 17.9400 0.7807 0.7772
nllb-back-syn-1:3 12.0352 0.7628 0.7493
opus-back-bl 18.1496 0.7811 0.7816
opus-back-syn-1:1-NLLB+ 17.9346 0.7781 0.7715
Table 3: Automatic Evaluation Scores from Different Models in CantonMT.
Hyperparameter Value
Learning Rate $10_{-4}$
Weight Decay 0.01
FP16 True
Table 4: Fine-tuning Hyperparameters w/ Hugging Face Trainer API